# PortLLM: Personalizing Evolving Large Language Models with Training-Free and Portable Model Patches

### v2

@@ 1 Introduction
The rise of large pretrained language models has marked a significant shift in natural language processing (NLP), enabling remarkable performance across a wide array of tasks, including natural language understanding and generation. These models, such as GPT-4 (Achiam et al., 2023), have achieved state-of-the-art performance by leveraging vast amounts of pretraining data and learning rich semantic representations that encode broad linguistic and world knowledge, which in turn allows them to generalize very well across domains (Antoniades et al., 2024). Despite their powerful general ability, pretrained LLMs often fall short in specialized domains where a nuanced understanding of technical language or context-specific knowledge is required (Wang et al., 2022a; Bommasani et al., 2021; Qiu et al., 2020). Hence, while pretraining enables LLMs to capture general knowledge, fine-tuning (e.g., personalization) is essential for adapting these models to specific domains. Fine-tuning bridges this gap by adapting pretrained models to specific tasks, enhancing their performance in domains such as healthcare, legal analysis, or scientific research (Min et al., 2021; Wei et al., 2021; Ouyang et al., 2022; Wang et al., 2022b; Liu et al., 2022; Raffel et al., 2020; Chen et al., 2024; Gao et al., 2024b). For instance, fine-tuned models can more effectively recognize domain-specific terminology, reason about complex relationships, and deliver more accurate and contextual appropriate responses.
@@
Much effort has been devoted to developing fine-tuning methods. Traditionally, fine-tuning would normally involve updating all the parameters of a model. For example, in the case of Mistral 7B (Jiang et al., 2023), it would involve updating all 7 billion parameters. Typically LLMs have billions of parameters, and so this process poses significant challenges in computational and memory requirements. To alleviate these constraints, many Parameter Efficient Fine-Tuning (PEFT) methods (Houlsby et al., 2019) have been proposed. One popular method is Low-Rank Adaptation (LoRA) (Hu et al., 2021), which aims to estimate an updated matrix $\Delta W$ using the product of two low-rank matrices $A$ and $B$. However, although LoRA lowers the training complexity, it still requires fine-tuning a large number of trainable parameters to reach a satisfactory performance. For example, LoRA fine-tuning Llama 2 13B (Touvron et al., 2023) variant would require up to eight A6000 GPUs with 48 GB VRAM each, for a very small batch size, hence imposing a considerable memory and computational burden.
@@
Furthermore, as cloud-hosted LLMs like ChatGPT (OpenAI, 2022) and Gemini (Team et al., 2023) undergo periodic (bi-annually or shorter) updates with newer data, their improved performance often renders previous versions outdated. For downstream users who have already invested in fine-tuning these models for domain-specific tasks, repeatedly fine-tuning or performing personalization at every new update is highly impractical, as the process is not only computationally expensive but also time-consuming.
@@
Beyond the computational and logistical hurdles, continual updates present another challenge for the downstream user: the lack of availability of the fine-tuning dataset. In domains such as healthcare or finance, data access is regulated by privacy laws and potentially time-sensitive. For example, fine-tuning medical models on patient data requires strict adherence to ethical and legal guidelines, making it difficult to repeatedly fine-tune with every model update. As a result, repeatedly fine-tuning models on newer LLM releases is not a viable long-term strategy for many users, hindering the downstream users from harnessing the performance gains from the evolving nature of LLMs. In response to these challenges, a nature research question comes:
@@
(Q) How to leverage the personalized knowledge captured in the first fine-tuned model to update any evolving LLMs?
@@
To address this question, we introduce PortLLM, training-free framework that enables seamless transfer of domain-specific knowledge across evolving models. PortLLM leverages model patches derived from LoRA, allowing users to port fine-tuned knowledge from one model iteration to another while preserving or even enhancing the performance of a downstream task. We show that our model patches are portable across model updates. If the downstream user fine-tunes a version of the model that has long become obsolete, they can simply add the model patches to the newer model, maintaining or boosting their performance on the downstream task. PortLLM eliminates the need for costly periodic fine-tuning, offering a scalable solution for maintaining task-specific performance across different model versions. Our contributions can be summarized as follows:
@@
  • We introduce PortLLM, a training-free framework designed to transfer knowledge between different versions of an evolving LLM. Given two model versions, PortLLM leverages task-specific model patches extracted from a pretrained LLM and seamlessly applies them to the updated version of the pretrained LLM. This process allows the updated model to achieve comparable, and in some cases improved, performance on downstream tasks — without any need for fine-tuning.
@@
  • Why do our model patches work? We address this question through both theoretical analysis and empirical investigation. Our findings reveal that certain terms in the model patch are effectively negligible, enabling us to create a simplified version of the patch that requires no training. Consequently, adding the simplified model patches alone is sufficient to achieve improved performance on the downstream task. Furthermore, we examine the impact of the pretraining dataset on downstream tasks, demonstrating that our framework can harness the benefits of continued pretraining across different model updates, across different pretraining datasets including {OpenOrca, SlimOrca, OpenPlatypus, AlpacaGPT4}.
@@ 2 Related Works
LLMs have transformed natural language processing, enabling models to perform complex tasks with remarkable accuracy and generalization. Models like GPT-3 (Brown et al., 2020), BERT (Devlin et al., 2019), and T5 (Raffel et al., 2023) have set benchmarks across a range of NLP tasks, from translation and summarization to question answering and text generation (Vaswani et al., 2023; Zhang et al., 2020; Rajpurkar et al., 2016). More recently, models like Llama (Touvron et al., 2023; Dubey et al., 2024), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) have pushed the boundaries further by optimizing both performance and computational efficiency. Llama was particularly notable for achieving state-of-the-art results with smaller model size, emphasizing that scaling down can still maintain performance. Mistral improves upon this by offering more efficient training methods, allowing faster training times without sacrificing model quality. Similarly, Gemma leverages architectural innovations to enhance model training and inference, focusing on reducing resource requirements while maintaining strong performance. Despite these advances, the computational demands for pretraining LLMs remain extremely high. Training large models requires weeks of computation on high-performance hardware, and continuous pretraining—to keep models updated with the latest data—adds further complexity and cost (Brown et al., 2020). Even after pretraining, the process of fine-tuning these models for specific downstream tasks remains resource-intensive, making the deployment of LLMs at scale a significant challenge (Raffel et al., 2023). This cost barrier continues to limit the accessibility and widespread adoption of LLMs across diverse applications. In this paper, we propose a training-free solution that enables the seamless transfer of personalized knowledge across evolving LLMs, reducing the need for costly fine-tuning and enhancing accessibility.
@@
The rapid growth in the size of pretrained LLMs has posed significant challenges for efficiently fine-tuning LLMs to specific downstream tasks. To address this, numerous PEFT methods have been developed, aiming to balance efficiency and accuracy. Early approaches focused on inserting trainable adapters—feed-forward networks placed between the layers of the pretrained model (Houlsby et al., 2019; Lin et al., 2020). Recent advancements have led to more sophisticated adapter-based PEFT methods (Mahabadi et al., 2021; Pfeiffer et al., 2020; Luo et al., 2023). A notable example is LoRA (Hu et al., 2021), which introduces trainable low-rank weight perturbations to the pretrained model, significantly reducing the number of parameters required for fine-tuning. LoRA’s key innovation lies in its use of the product of two low-rank matrices to approximate weight changes. Building upon this concept, Q-LoRA (Dettmers et al., 2023) proposes to propagate gradients through LoRA using a $4$-bit quantized pretrained model, further enhancing efficiency. Concurrently, prompt-based learning methods have demonstrated effectiveness across various NLP tasks. Methods such as prompt-tuning (Lester et al., 2021) and prefix-tuning (Li & Liang, 2021) incorporate learnable continuous embeddings into the model’s hidden states. They condition the frozen model to adapt to specific tasks without modifying the underlying architecture. Despite these advances, fine-tuning each updated LLM with PEFT to equip personalized knowledge remains highly costly, and how PEFT can bridge the gap in personalized settings within this evolving environment in a portable manner is yet to be fully explored. To this end, we develop in this paper the theory behind portable model patches that can be plugged into an updated model to carry over the personalized knowledge from the first fine-tuned model.
3 Methodology
@@ 3.1 Preliminaries
Figure 1: The diagram illustrates the core components of PortLLM, a training-free framework to port personalized knowledge between evolving LLMs. Initially, a pretrained LLM is fine-tuned using LoRA. We transfer this task-specifc knowledge without requiring the newer updated model to be fine-tuned again. This allows for continual performance improvements on downstream tasks without additional periodic fine-tuning.
@@
Transformer models (Vaswani et al., 2023) is an architecture that has revolutionized NLP and other sequence-based tasks. It consists of two key components: (1) Multi-Head Self-Attention and (2) Feed-Forward Neural Network. The Multi-Head Self-Attention mechanism is the core innovation of transformers. It computes the weighted representation of the input sequence where each token attends to every other token. Suppose we are given an input $X\in\mathbb{R}^{n\times d}$ where $n$ refers to the sequence length and $d$ is the hidden dimension of the transformer model. Then for any given attention head $i$ with a total of $H$ heads, we define the following three matrices: query matrix $W_{q_{i}}\in\mathbb{R}^{d\times d_{H}}$, key matrix $W_{k_{i}}\in\mathbb{R}^{d\times d_{H}}$ and value matrix $W_{v_{i}}\in\mathbb{R}^{d\times d_{H}}$, where $d_{H}=d/H$. Given these, we can compute the self-attention for the $i$th head as follows
@@
where $W_{o}\in\mathbb{R}^{d\times d}$ is a projection matrix that combines the outputs of the different heads back into the model’s hidden dimension $d$. After the Muli-head Self-Attention, the output is passed through a position-wise feed-forward network that consists of two linear transformations and a nonlinear activation function (like ReLU or GELU). Given the weight matrix of the first linear layer $W_{\text{up}}\in\mathbb{R}^{d\times d_{m}}$, the weight matrix of the second layer $W_{\text{down}}\in\mathbb{R}^{d_{m}\times d}$, bias terms $b_{1}\in\mathbb{R}^{d_{m}}$, $b_{2}\in\mathbb{R}^{d}$, and a non-linear activation function $\sigma(\cdot)$ where $d_{m}$ is the hidden dimension, Feed-Forward Network is applied independently to each position in the sequence as follows
@@
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times d}$ are trainable matrices, while $r$ is the rank of the decomposition with $r\ll d$. In this setup, the full weight update matrix $\Delta W$ is replaced by the product of two smaller matrices, $B$ and $A$, drastically reducing the number of trainable parameters from $d^{2}$ to $2rd$. Hence, the fine-tuned model weights we can be trained as follows:
@@ 3.3 Analysis of Our Proposed Portability
The fine-tuned updated model $\theta^{\prime}_{i}$ can be decomposed into a naive update term and a residual matrix $C$ as follows:
@@
$=\underbrace{(\theta^{\prime}+\Delta\theta_{i})}_{\text{Naive Update}}+\underbrace{(\Delta\theta_{i}^{\prime}-\Delta\theta_{i})}_{\text{Residual Matrix }C}.$ (7b)
Lemma: We claim that the residual matrix is negligible compared to the naive update. Proof: First, recall the assumption in Section 3.2 that $\Delta\theta_{i}^{\prime}$ and $\Delta\theta_{i}$ have ranks that are much smaller than the rank of $\Delta\theta$. Second, $\theta^{\prime}=\theta+\Delta\theta$ ensures that $\operatorname{rank}(\theta^{\prime})\geq\operatorname{rank}(\theta)$. We therefore conclude that the rank of the residual matrix is approximately much smaller than that of the naive update matrix.
@@
We empirically show that the difference between two personalization updates, $C=\Delta\theta_{i}^{\prime}-\Delta\theta_{i}$, is negligible when compared to naive update term, $\theta^{\prime}+\Delta\theta_{i}$. We use Frobenius norm, $\|\cdot\|_{F}$, and the maximum singular value, $\sigma_{\max}$, to measure the magnitudes of matrices. The results are summarized in Table 1 across four different downstream tasks {BoolQ, MRPC, RTE, WNLI}. We can see that the $\sigma_{\max}$ for first term are on average $\times6$ and the $\|\cdot\|_{F}$ is 136×136$\times$136 × bigger in the favor of the first term, which implies that $C$ is comparatively negligible, hence our model patch can be simplified as shown in (6).
@@ 4 Experiments
We evaluate our framework on a diverse set of datasets to demonstrate PortLLM universality and effectiveness across various downstream tasks and domains. Specifically, we leverage datasets from the GLUE (Wang, 2018), SuperGLUE (Wang et al., 2019), WinoGrande (Sakaguchi et al., 2021) and GSM8K (Cobbe et al., 2021) benchmarks commonly used for such evaluation in literature. For Question Answering tasks, we utilize BoolQ (from SuperGLUE) and SST-2 (from GLUE); for Similarity and Paraphrase Tasks, the MRPC (from GLUE) dataset; for Inference Tasks, the RTE and WNLI (both from GLUE) datasets; and lastly for Reasoning Tasks, we employ WinoGrande and GSM8K. This broad spectrum of tasks enables a comprehensive evaluation of our model’s performance across diverse downstream applications. For continued pretraining datasets, we use the following: OpenOrca (Lian et al., 2023a), SlimOrca (Lian et al., 2023b), OpenPlatypus (Lee et al., 2023) and AlpacaGPT4 (Peng et al., 2023). Additionally, we conduct extensive experiments across multiple model architectures to demonstrate the generalizablity of our framework. Specifically, we test our method on Mistral-7B (Jiang et al., 2023), Llama2-7B (Touvron et al., 2023), Llama3.1-8B (Dubey et al., 2024), and Gemma2-9B (Team et al., 2024), showcasing the robustness and adaptability of our approach across different LLMs.
@@
To simulate the progression of time, we employ continued pretraining, where we transition from $\theta$ to $\theta^{\prime}$ by taking a pretrained model and further pretraining it on a specific dataset using LoRA (Hu et al., 2021). In all continued pretraining scenarios, we maintain a constant rank $r=64$ and $\alpha=128$ with a learning rate of $0.0001$ and $4$ epochs. For downstream tasks, we also use LoRA, but in this case, we set the rank $r=8$ consistently to ensure a fair comparison across tasks. Furthermore, for all LoRA applications we optimize all the attention layers (Key, Value, Query, Projection), and all Feed Forward Network layers (Up Projection, Down Projection and Gate Projection, where applicable). For each downstream fine-tuning, we use a constant learning rate of $0.0004$ while the number of epochs for each dataset {BoolQ, SST-2, MRPC, RTE, WinoGrande, WNLI, GSM8K} is as follows {$5,5,5,5,3,5,1$}. Lastly, for fine-tuning on specific downstream dataset mentioned, we solely use the train split, and evaluate on test split.
@@
We use the Language Model Evaluation Harness (Gao et al., 2024a) by EleutherAI to assess the performance of all trained and fine-tuned models across the datasets in our experiments. All evaluations are conducted in a zero-shot setting rather than a few-shot setting. For datasets {BoolQ, SST-2, RTE, WinoGrande, WNLI}, we employ accuracy as the primary evaluation metric. In the case of {MRPC}, we utilize both accuracy and F1 score to provide a more comprehensive evaluation. For {GSM8K}, we had two evaluation options: (1) Flexible Match Accuracy or (2) Exact Match Accuracy. Due to poor zero-shot performance of the models on GSM8K for exact matches, we opted for Flexible Match Accuracy.

### v3

@@ 1 Introduction
The rise of pretrained large language models (LLMs) has marked a significant paradigm shift in natural language processing (NLP), particularly in their ability to adapt to specific domains and tasks. These models, such as GPT-4 (Achiam et al., 2023), have achieved state-of-the-art performance by leveraging vast amounts of pretraining data (Antoniades et al., 2024). However, pretrained LLMs often require adaptation for specialized domains where context-specific knowledge is critical  (Wang et al., 2022a; Bommasani et al., 2021; Qiu et al., 2020). Hence, while pretraining provides a strong foundation, fine-tuning (e.g., personalization) is essential for specific domains. Fine-tuning bridges this gap by adapting pretrained models to specific tasks, enhancing their performance in domains such as healthcare, legal analysis, or scientific research (Min et al., 2021; Wei et al., 2021; Ouyang et al., 2022; Wang et al., 2022b; Liu et al., 2022; Raffel et al., 2020; Chen et al., 2024; Gao et al., 2024b). For instance, fine-tuned models can more effectively recognize domain-specific terminology, reason about complex relationships, and deliver more accurate and contextual appropriate responses.
@@
Much effort has been devoted to developing fine-tuning methods. Traditionally, fine-tuning would normally involve updating all the parameters of a model. For example, in the case of Mistral 7B (Jiang et al., 2023), it would involve updating all 7 billion parameters. Typically, LLMs have billions of parameters, and so this process poses significant challenges in computational and memory requirements. To alleviate these constraints, many parameter efficient fine-tuning (PEFT) methods (Houlsby et al., 2019) have been proposed. One popular method is low-rank adaptation (LoRA) (Hu et al., 2021), which aims to estimate an update matrix $\Delta W$ using the product of two low-rank matrices $A$ and $B$. However, although LoRA lowers the training complexity, it still requires fine-tuning a large number of trainable parameters to reach a satisfactory performance. For example, LoRA fine-tuning a Llama 2 13B (Touvron et al., 2023) variant would require up to eight A6000 GPUs with 48 GB VRAM each, for a very small batch size, hence imposing a considerable memory and computational burden.
@@
Furthermore, as cloud-hosted LLMs like ChatGPT (OpenAI, 2022) and Gemini (Team et al., 2023) undergo periodic (biannual or more frequent) updates with newer data, the best performing model often renders previous versions outdated. For downstream users who have already invested in fine-tuning the previous models for domain-specific tasks, repeatedly fine-tuning or performing personalization at every new update is highly impractical, as this process is not only computationally expensive but also time-consuming.
@@
Beyond the computational and logistical hurdles, continual updates present another challenge for the downstream user: the lack of availability of the fine-tuning dataset. In domains such as healthcare or finance, data access is regulated by privacy laws and potentially time-sensitive. For example, fine-tuning medical models on patient data requires strict adherence to ethical and legal guidelines, making it difficult to repeatedly fine-tune with every LLM release. As a result, repeatedly fine-tuning models on newer LLM releases is not a viable long-term strategy for many users, hindering the downstream users from harnessing the performance gains from the evolving nature of LLMs. In response to these challenges, a natural research question arises:
@@
(Q) How to leverage the personalized knowledge captured in the first fine-tuned LLM model to update any evolving LLMs?
@@
To address this question, we introduce PortLLM, a training-free framework that enables seamless transfer of domain-specific knowledge across evolving models. PortLLM leverages model patches derived from LoRA, allowing users to port fine-tuned knowledge from one model iteration to another while preserving or even enhancing the performance of a downstream task. We show that our model patches are portable across model updates. If the downstream user fine-tuned a version of the model that has long become obsolete, they can simply add the model patch to the newer model, maintaining or boosting their performance on the downstream task. PortLLM eliminates the need for costly periodic fine-tuning, offering a scalable solution for maintaining task-specific performance across different model versions. Our contributions can be summarized as follows:
@@
  • We introduce PortLLM, a training-free framework designed to transfer knowledge between different versions of an evolving LLM. Given two LLM versions, PortLLM leverages task-specific model patches extracted from a fine-tuned LLM and seamlessly applies them to the evolved LLM. This process allows the updated LLM to achieve comparable, and in some cases improved, performance on downstream tasks—without any need for fine-tuning.
@@
  • Why do our model patches work? We address this question through both theoretical analysis and empirical investigation. Our findings reveal that certain terms in the model patch are effectively negligible, enabling us to create a simplified version of the patch that requires no training. Consequently, adding the simplified model patches alone is sufficient to achieve improved performance on the downstream task. Furthermore, we examine the impact of the pretraining dataset on downstream tasks, demonstrating that our framework can harness the benefits of continued pretraining across different model updates and across different pretraining datasets including {OpenOrca, SlimOrca, OpenPlatypus, AlpacaGPT4}.
@@ 2 Related Works
LLMs have transformed natural language processing, enabling models to perform complex tasks with remarkable accuracy and generalization. Models like GPT-3 (Brown et al., 2020), BERT (Devlin et al., 2019), and T5 (Raffel et al., 2023) have set benchmarks across a range of NLP tasks, from translation and summarization to question answering and text generation (Vaswani et al., 2023; Zhang et al., 2020; Rajpurkar et al., 2016). More recently, models like Llama (Touvron et al., 2023; Dubey et al., 2024), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) have pushed the boundaries further by optimizing both performance and computational efficiency. LLama, Mistral, and Gemma represent recent advances in LLM architectures, each offering improvements in efficiency and performance. However, even with such improvements, the performance on domain-specific downstream tasks is subpar, making fine-tuning necessary. In this paper, we propose a training-free solution that enables the seamless transfer of personalized knowledge across evolving LLMs, reducing the need for costly fine-tuning and enhancing accessibility.
@@
The rapid growth in the size of pretrained LLMs has posed significant challenges for efficiently fine-tuning LLMs to specific downstream tasks. To address this challenge, numerous PEFT methods have been developed, aiming to balance efficiency and accuracy. Early approaches focused on inserting trainable adapters—feed-forward networks placed between layers of the pretrained model (Houlsby et al., 2019; Lin et al., 2020). Recent advancements have led to more sophisticated adapter-based PEFT methods (Mahabadi et al., 2021; Pfeiffer et al., 2020; Luo et al., 2023) including LMaaS (Sun et al., 2022) for service-oriented adaptation and kNN-Adapter (Huang et al., 2023) for retrieval-augmented fine-tuning. A notable example is LoRA (Hu et al., 2021), which introduces trainable low-rank weight perturbations to the pretrained model, significantly reducing the number of parameters required for fine-tuning. LoRA’s key innovation lies in its use of the product of two low-rank matrices to approximate weight changes. Building upon this concept, several methods have emerged including Q-LoRA (Dettmers et al., 2023), CombLM (Ormazabal et al., 2023), and IPA (Lu et al., 2023). Concurrently, prompt-based learning methods have demonstrated effectiveness across various NLP tasks. Methods such as prompt-tuning (Lester et al., 2021), prefix-tuning (Li & Liang, 2021) and more recent approaches like proxy-tuning (Liu et al., 2024) and BBox-Adadpter (Sun et al., 2024) incorporate learnable continuous embeddings into the model’s hidden states. They condition the frozen model to adapt to specific tasks without modifying the underlying architecture. Despite these advances, fine-tuning each updated LLM with PEFT to equip personalized knowledge remains highly costly, and how PEFT can bridge the gap in personalized settings within this evolving environment in a portable manner is yet to be fully explored. To this end, we develop in this paper the theory behind portable model patches that can be plugged into an evolved LLM to carry over the personalized knowledge from an earlier fine-tuned LLM.
3 Proposed Method
@@ 3.1 Preliminaries
Figure 1: The diagram illustrates the core components of PortLLM, a training-free framework to port personalized knowledge between evolving LLMs. Initially, a pretrained LLM is fine-tuned using LoRA. We transfer this task-specific knowledge without requiring the newer updated model to be fine-tuned again. This allows for continual performance improvements on downstream tasks without additional fine-tuning.
@@
Transformer models (Vaswani et al., 2023) is an architecture that has revolutionized NLP and other sequence-based tasks. It consists of two key components: (1) Multi-head self-attention and (2) feed-forward network. The multi-head self-attention mechanism is the core innovation of transformers. It computes the weighted representation of the input sequence where each token attends to every other token. Suppose we are given an input $X\in\mathbb{R}^{n\times d}$ where $n$ is the sequence length and $d$ is the hidden dimension of the transformer model. Then for any given attention head $i$ with a total of $H$ heads, we define the following three matrices: query matrix $W_{q_{i}}\in\mathbb{R}^{d\times d_{H}}$, key matrix $W_{k_{i}}\in\mathbb{R}^{d\times d_{H}}$, and value matrix $W_{v_{i}}\in\mathbb{R}^{d\times d_{H}}$, where $d_{H}=d/H$. Given these, we can compute the self-attention for the $i$th head as follows:
@@
where $W_{o}\in\mathbb{R}^{d\times d}$ is a projection matrix that combines the outputs of the different heads back into the model’s hidden dimension $d$. After the multi-head self-attention, the output is passed through a position-wise feed-forward network that consists of two linear transformations and a nonlinear activation function (e.g., ReLU or GELU). Given the weight matrix of the first linear layer $W_{\text{up}}\in\mathbb{R}^{d\times d_{m}}$, weight matrix of the second layer $W_{\text{down}}\in\mathbb{R}^{d_{m}\times d}$, bias terms $b_{1}\in\mathbb{R}^{d_{m}}$, $b_{2}\in\mathbb{R}^{d}$, and a nonlinear activation function $\sigma(\cdot)$, where $d_{m}$ is the hidden dimension, the feed-forward network is applied independently to each position in the sequence as follows
@@
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times d}$ are trainable matrices, while $r$ is the rank of the decomposition with $r\ll d$. In this setup, the full weight update matrix $\Delta W$ is replaced by the product of two smaller matrices, $B$ and $A$, drastically reducing the number of trainable parameters from $d^{2}$ to $2rd$. Hence, the fine-tuned model weights can be trained as follows:
@@ 3.3 Analysis of Our Proposed Portability
The fine-tuned updated model $\theta^{\prime}_{i}$ can be decomposed into a naive update term and a residual matrix $R$ as follows:
@@
$=\underbrace{(\theta^{\prime}+\Delta\theta_{i})}_{\text{Naive Update }\hat{\theta}_{i}^{\prime}}+\underbrace{(\Delta\theta_{i}^{\prime}-\Delta\theta_{i})}_{\text{Residual Matrix }R}.$ (7b)
Lemma 1 (informal): We claim that the residual matrix $R$ is negligible compared to the naive update $\hat{\theta}_{i}^{\prime}$. The key reason is that the model patches $\Delta\theta_{i}^{\prime}$ and $\Delta\theta_{i}$ are both low rank (recall the assumption in Section 3.2), whereas the pretrained models $\theta$ and $\theta^{\prime}$ are primarily full rank matrices. We provide a formal proof in Appendix C.
@@
We empirically show that the difference between two personalization updates, $R=\Delta\theta_{i}^{\prime}-\Delta\theta_{i}$, is negligible when compared to the naive update term, $\theta^{\prime}+\Delta\theta_{i}$. We use Frobenius norm, $\|\cdot\|_{F}$, and the maximum singular value, $\sigma_{\max}$, to measure the magnitudes of matrices. The results are summarized in Table 1 across four different downstream tasks {BoolQ, MRPC, RTE, WNLI}. We can see that the $\sigma_{\max}$ for first term are on average $\times6$ and the $\|\cdot\|_{F}$ is 136×136$\times$136 × bigger in the favor of the first term, which implies that $R$ is comparatively negligible, hence our model patch can be simplified as shown in (6).
@@ 4 Experiments
We evaluate our framework on a diverse set of datasets to demonstrate PortLLM’s universality and effectiveness across various downstream tasks and domains. Specifically, we leverage datasets from the GLUE (Wang, 2018), SuperGLUE (Wang et al., 2019), WinoGrande (Sakaguchi et al., 2021) and GSM8K (Cobbe et al., 2021) benchmarks commonly used for such evaluation in the literature. For question answering tasks, we utilize BoolQ (from SuperGLUE) and SST-2 (from GLUE); for similarity and paraphrase tasks, the MRPC (from GLUE) dataset; for inference tasks, the RTE and WNLI (both from GLUE) datasets; and lastly for reasoning tasks, we employ WinoGrande and GSM8K. This broad spectrum of tasks enables a comprehensive evaluation of our model’s performance across diverse downstream applications. For continued pretraining datasets, we use the following: OpenOrca (Lian et al., 2023a), SlimOrca (Lian et al., 2023b), OpenPlatypus (Lee et al., 2023) and AlpacaGPT4 (Peng et al., 2023). Additionally, we conduct extensive experiments across multiple model architectures to demonstrate the generalizability of our framework. Specifically, we test our method on Mistral-7B (Jiang et al., 2023), Llama2-7B (Touvron et al., 2023), Llama3.1-8B (Dubey et al., 2024), and Gemma2-9B (Team et al., 2024), showcasing the robustness and adaptability of our approach across different LLMs.
@@
To simulate the progression of time, we employ continued pretraining, where we transition from $\theta$ to $\theta^{\prime}$ by taking a pretrained model and further pretraining it on a specific dataset using LoRA (Hu et al., 2021). In all continued pretraining scenarios, we maintain a constant rank $r=64$ and $\alpha=128$ with a learning rate of $0.0001$ and $4$ epochs. For downstream tasks, we also use LoRA, but in this case, we set the rank $r=8$ consistently to ensure a fair comparison across tasks. Furthermore, for all LoRA applications, we optimize all the attention layers (Key, Value, Query, Projection) and all feed-forward network layers (Up Projection, Down Projection, and Gate Projection, where applicable). For each downstream fine-tuning, we use a constant learning rate of $0.0004$ while the number of epochs for each dataset {BoolQ, SST-2, MRPC, RTE, WinoGrande, WNLI, GSM8K} is {$5,5,5,5,3,5,1$}, respectively. Lastly, for fine-tuning on a specific downstream dataset mentioned, we solely use the train split and evaluate on the test split.
@@
We use the Language Model Evaluation Harness (Gao et al., 2024a) by EleutherAI to assess the performance of all trained and fine-tuned models across the datasets in our experiments. All evaluations are conducted in a zero-shot setting rather than a few-shot setting. For datasets {BoolQ, SST-2, RTE, WinoGrande, WNLI}, we employ accuracy as the primary evaluation metric. In the case of {MRPC}, we utilize both accuracy and F1 score to provide a more comprehensive evaluation. For {GSM8K}, we had two evaluation options: (1) flexible match accuracy or (2) exact match accuracy. Due to the poor zero-shot performance of the models on GSM8K for exact matches, we opted for the flexible match accuracy.